Hello. Please help with this.
I have to write the equation of y+ x^2 + 6x + 5 in vertex form. Showing all work.
I appreciate your help!

Answer by stanbon(57307)

y = x^2 + 6x + 5 in vertex form
---
y = (x^2+6x+?) + 5-?
-----
Complete the square:
y = (x^2+6x+9) + 4 - 9
---
y = (x+3)^2-5
---------------------
Cheers,
Stan H.